OpenStudy (anonymous):

Find derivative of (60x^(-4/3))

OpenStudy (anonymous):

\[\frac{dy}{dx}=(60{x}^{-\frac{4}{3}})'=60*(-\frac{4}{3}){x}^{-\frac{4}{3}-1}=-80{x}^{-\frac{7}{3}}\]